OpenStudy (christos):

integral random question, Can I split this integral into one integral of "2" minus another integral of the "fraction" ?

OpenStudy (christos):

Ops forgot to mention the integral : http://screencast.com/t/sFjYCiQ4R

OpenStudy (christos):

There it is* ^

hartnn (hartnn):

yes, you can`

OpenStudy (christos):

ok sec

OpenStudy (christos):

so my answer to this integral is 2x - 16xln(x^2 + 4) + c . Is it correct ?

hartnn (hartnn):

nopes.

hartnn (hartnn):

\(\Large \int \dfrac{1}{x^2+a^2}dx=\dfrac{1}{a}\tan^{-1}({\dfrac{x}{a}})\)

hartnn (hartnn):

here a^2= 4

OpenStudy (dls):

yup.

OpenStudy (dls):

\[\LARGE 2 \int\limits dx-8 \int\limits \frac{dx}{x^2+4}\]

OpenStudy (christos):

I know this way at @hartnn but I just tried to do it another way without involving reverse functions, you understand what I was trying to do ?

hartnn (hartnn):

no way to do it without tan inverse.

hartnn (hartnn):

to involve ln(x^2+4), there should have been 2x in the numerator, but there isn't.

OpenStudy (christos):

I see, thanks!

hartnn (hartnn):

welcome ^_^